# DynGMA: a robust approach for learning stochastic differential equations from data

### v1

@@ 4.1 Improving accuracy of density approximation
Here we consider three kinds of approximations for the exact density: (i) the one-step Euler-Maruyama approach, i.e., Gaussian approximation with mean $x_{0}+t\tanh x_{0}$ and covariance $\sqrt{t}$; (ii) the one-step Gaussian approximation , i.e., numerically solving Equation (13) using (14) with $L=2$ to obtain the mean and covariance; and (iii) the proposed DynGMA using Algorithm 1 with $h=0.5$ and $L=2$.
With an initial condition of $x_{0}=1/2$, the comparison of these approximations to the exact density is illustrated in Fig. 1. As depicted in the figure, the proposed DynGMA demonstrates a significantly more accurate approximation, especially noticeable when the time step is large. Furthermore, considering that the one-step Euler-Maruyama scheme is inherently a Gaussian approximation, the leftmost figure highlights that the numerical solution of Equation (13) yields more accurate mean and covariance results. These findings demonstrate that numerically solving Equation (13) with utilizing (14), instead of using the Euler-Maruyama, in each substep to obtain the Gaussian approximation within the proposed DynGMA framework, facilitates a more accurate approximation of the transition density.
@@ 4.2 Improving accuracy of learning SDEs
In this subsection, we demonstrate that, compared to the Euler-Maruyama approach, learning SDEs using DynGMA (i) exhibits superior accuracy with a given step size; (ii) can handle larger step sizes effectively; and (iii) performs well even in the presence of higher levels of measurement noise. As the first illustration of learning SDEs, we examine a two-dimensional system that has been studied for computing invariant distributions from data lin2023computing . In that research, the Euler-Maruyama approach is employed to learn the governing functions, and the invariant distribution is computed simultaneously. Here, we use the proposed DynGMA to reconstruct the entire dynamics and then compute the invariant distribution using their decomposition method. The system is formulated as:
@@
We generated $2000\left(\frac{\Delta t}{0.05}\right)$ trajectories for our training data. Each trajectory consists of an initial state uniformly sampled from $[-2,2]\times[-3,3]$, and includes $M=1/\Delta t$ states at equidistant time steps of $\Delta t$. In other words, there are a total of $4\times 10^{4}$ data points in the dataset. We employed an FNN with two hidden layers, each having 128 units, and using tanh activation function to parameterize the drift $f_{\theta}$. Additionally, We use a constant $2$-by-$2$ matrix to parameterize the diffusion $\sigma_{\theta}$. We use Algorithm 1 with $h=0.1$ and a step size of $0.05$ in (14) to approximate the transition density. During training, we employed full-batch Adam optimization for $5\times 10^{3}$ epochs, with the learning rate set to exponentially decay with linearly decreasing powers from $10^{-2}$ to $10^{-4}$.
@@
Figure 3: Errors in the learned diffusion coefficients and the generalized potential of the learned dynamics when using data with measurement noise for the two-dimensional system (19). The results are obtained by taking the mean of $5$ independent experiments, and the black error bar represents one standard deviation.
After completing the training process, we compute the invariant distributions and the generalized potential of the learned dynamics using the decomposition method lin2023computing , of which details are present in B. Subsequently, we plot the drift functions, the potential of the learned, and the exact dynamics in Fig. 2. Here, the reference solution is derived by numerically solving the FPK equation using the finite difference method. The results demonstrate agreements between both solutions when the step size $\Delta t$ is small. However, the Euler-Maruyama approach exhibits limitations in learning accurate dynamics and generalized potentials for larger step sizes. In contrast, DynGMA consistently provides accurate results, as evidenced in the third column of Fig. 2.
@@
Finally, we show that multi-step training using DynGMA improves robustness in the presence of measurement noise in the system observations. We generate $4000$ trajectories and sample $41$ states at equidistant time steps with $\Delta t=0.025$. Each trajectory point is multiplied by $1+\delta$ to generate the training dataset, where $\delta$ follows a uniform distribution on $[-\text{Noise},\text{Noise}]$, and the intensity Noise is set to $0.01$, $0.02$, and $0.04$. We optimize the multi-step loss (18), where $\Gamma=\{j\}$ and $j$ takes values $1,2,4,8$. As the Euler-Maruyama approach is only valid for one step, we directly expand its step size, take
@@
and substitute it into loss (18). Note that excessive measurement noise makes it challenging to distinguish between measurement noise and the true randomness induced by diffusion. Thus, we record the error of the learned diffusion and the learned potential in Fig. 3. It depicts that the error increases with the increase in noise, especially when using one-step training ($j=1$). Nevertheless, with multi-step training involving more steps, the impact of noise on the results gradually decreases for the proposed DynGMA. This trend is also noticeable in the Euler-Maruyama approach. However, increasing the number of steps in the Euler-Maruyama approach (20) results in a larger time step size, leading to significant numerical errors caused by discretization. These results demonstrate the effectiveness of DynGMA in handling data with measurement noise.
@@ 4.3 Learning SDEs with state-dependent diffusions
Subsequently, we consider the $3$d Lorenz system lorenz1963deterministic ; sparrow2012lorenz , which is a simplified model for atmospheric dynamics and is widely investigated for learning dynamics lin2023computing ; yu2021onsagernet . To showcase the effectiveness of the proposed method in learning SDEs with state-dependent diffusions, we investigate the Lorenz system in the presence of multiplicative noise chekroun2011stochastic , formulated as:
@@
We simulate $N=10^{4}$ trajectories for the training dataset. For each trajectory, the initial state is sampled from the uniform distribution on the domain $[-25,25]\times[-30,30]\times[-10,60]$, and we collect $M=200$ pairs of data points at the time slices $\{t_{j}=0.01*j:\ j=0,1,\cdots,M-1\}$. In total, there are $2\times 10^{6}$ pairs of points. Here, Algorithm 2 with $h=0.005$ is used to obtain the approximation of the transition density. And the results are compared with those obtained using the Euler-Maruyama and Gaussian Cubature methods (employing the Euler method with a step size of $0.005$). In all methods, the drift $f_{\theta}$ is parameterized in the same manner as described in Section 4.2, and the diffusion $\sigma_{\theta}$ is parameterized according to (16), where the used FNN has one hidden layer with $50$ units and tanh activations. The optimization process employs the Adam optimization algorithm, with a chosen batch size of $10^{5}$. The learning rate is set to $10^{-3}$ for DynGMA and the Euler-Maruyama and $10^{-4}$ for the Gaussian Cubature approach. Results are collected after $10^{5}$ parameter updates.
@@
Moreover, we compute the errors $e_{f}$, $e_{\sigma}$, and $e_{p}$, evaluated at $10^{5}$ points uniformly sampled in $[-25,25]\times[-30,30]\times[-10,60]$. The results are presented in Table 3 for various methods. It is shown that our proposed DynGMA outperforms other methods in terms of accuracy. Interestingly, despite the Gaussian Cubature having a discrete step size smaller than that of the Euler-Maruyama approach, it fails to yield a substantial improvement. It is noted that the Gaussian Cubature method involves two steps, and an excessive number of steps can potentially lead to non-positive covariance, resulting in the failure of the training process. Currently, the application of the Gaussian Cubature method in machine learning is limited to time series prediction using SDE as black-box solin2021scalable . Its utilization for learning interpretable drift and diffusion functions in SDE models has not been explored. Further research might be required to investigate the parameterization and discretization of this method for such tasks.
@@ 4.4 Learning high-dimensional SDEs
We next validate the performance of the DynGMA in learning a $10$d biological model combined by the epithelial-mesenchymal transition (EMT) network and the metastasis network. Researchers can investigate the correlation between EMT and the metastasis of cancer via this biological model heerboth2015emt and have proposed several works to compute its invariant distribution li2018landscape ; lin2023computing . The Euler-Maruyama loss has been employed to learn its stochastic dynamics and invariant distributions from data lin2023computing . Herein, we use a similar experimental setup but a larger time step size.
Specifically, the training data is obtained by solving the governing SDE with constant diffusion, of which the governing function is given in A. We first sample $10^{5}$ initial states from the uniform distribution on the domain $\Omega_{1}\times\cdots\times\Omega_{10}$ where $\Omega_{i}=[0,2]$ for $i\neq 7$ and $\Omega_{7}=[0,6]$, then simulate trajectories and collect $36$ pairs of data points at the times $\{t_{j},t_{j}+\Delta t\}$, where $t_{j}=2,2.5,\cdots,19.5$. In total, there are $7.2\times 10^{6}$ data points. The step size $\Delta t$ is set as $0.005$ in previous work lin2023computing and here we configure $\Delta t$ as $0.04$ or $0.08$. Due to considerations of computational efficiency and resources, we apply the Adam optimization with a batch size of $10^{5}$ for $10^{5}$ iterations to train the model, and we use Algorithm 1 with $K=1$ and $L=2$, i.e., Gaussian density approximation with (14). It is noted that in high-dimensional scenarios, DynGMA with larger $K$ would significantly increase computational complexity. Therefore, herein we solely employ the proposed Gaussian approximation. The chosen model architecture and other hyperparameters remain consistent with those outlined in Section 4.2.
@@ 4.5 Learning effective SDEs from Gillespie’s stochastic simulations data with variable time step-sizes
In the subsequent sections, we validate the effectiveness of the proposed DynGMA in handling data with variable, and even uncontrollable, time step-sizes. We consider learning effective SDEs from trajectories generated by Gillespie’s stochastic simulation algorithm (SSA), where an exact SDE does not exist. SSA generates sample sequences of continuous-time Markov chains, where individual events occur with random timing. The time steps between reaction events follow an exponential distribution, resulting in variable and uncontrollable step sizes. Specifically, We consider the Susceptible-Infectious-Recovered (SIR) model, an epidemiological compartment model used to simulate the spread of infectious diseases in a population. In this model, individuals can transition between three states: Susceptible (S), Infectious (I), and Recovered (R). Let $y_{1}$ and $y_{2}$ denote the concentrations of infected and recovered individuals, respectively. By performing Gillespie’s SSA for the SIR model, we obtain the stochastic dynamics of $y_{1}$ and $y_{2}$. The mass balance then yields that the concentration of susceptible individuals is given by $y_{0}=1-y_{1}-y_{2}$.
@@
In Gillespie’s SSA, the parameters $t_{max}$ and $\Delta t$ are predetermined, and the data points are evaluated at ${t_{0},t_{1},\cdots,t_{max}}$, where $\Delta t_{i}=t_{i}-t_{i-1}\approx\Delta t$. In this experiment, we set $t_{max}=1$ and choose various values for $\Delta t$ to generate $10^{5}$ trajectories for each training dataset and $2.5\times 10^{4}$ trajectories for each test dataset. The data step size is a random variable and is illustrated in Fig. 6 for different settings of $\Delta t$. There exist some uncontrollably large steps $\Delta t_{i}$ for all settings of $\Delta t$. We use one hidden layer in FNN with $64$ units and tanh activations to represent drift $f_{\theta}$, and parameterize diffusion $\sigma_{\theta}$ according to (16). Adam optimization with the learning rate of $10^{-3}$ is used to update parameters for $2\times 10^{4}$ epochs. Under these settings, we carry out Algorithm 2 (if $\Delta t=0.2$, then $L=2$; otherwise, $L=1$) to obtain the approximation and optimize the loss function (17) to train the networks.
@@
Next, we consider the SIRS model, where recovered individuals have a probability of being reinfected and richer dynamic behaviors are exhibited. We generate the data for the SIRS model in the same manner and use the same settings to train networks. The predictions are shown in Fig. 9. It is shown that the learned network model captures stochastic dynamics perfectly and is capable of reconstructing the correct oscillations and steady state.

### v2

@@ 4.1 Improving accuracy of density approximation
Here we consider three kinds of approximations for the exact density: (i) the one-step Euler-Maruyama approach, i.e., Gaussian approximation with mean $x_{0}+t\tanh x_{0}$ and covariance $\sqrt{t}$; (ii) the one-step Gaussian approximation , i.e., numerically solving Equation (13) using (14) with $L=2$ to obtain the mean and covariance; and (iii) the proposed DynGMA using Algorithm 1 with parameter settings specified as $h=0.5$ corresponding to $K=t/0.5$ and $L=2$.
With an initial condition of $x_{0}=1/2$, the comparison of these approximations to the exact density is illustrated in Fig. 2. As depicted in the figure, the proposed DynGMA demonstrates a significantly more accurate approximation, especially noticeable when the time step is large. Furthermore, considering that the one-step Euler-Maruyama scheme is inherently a Gaussian approximation, the leftmost figure highlights that the proposed Gaussian approximation yields more accurate mean and covariance results. These findings demonstrate that numerically solving Equation (13) with utilizing (14), instead of using the Euler-Maruyama, in each sub-step to obtain the Gaussian approximation within the proposed DynGMA framework, facilitates a more accurate approximation of the transition density. Furthermore, the proposed DynGMA is also capable of learning more accurate approximation of the drift and diffusion functions from data. Experimental results and details are presented in C.
@@ 4.2 Improving accuracy of learning SDEs
In this subsection, we demonstrate that, compared to the Euler-Maruyama approach, learning SDEs using DynGMA (i) exhibits superior accuracy with a given step size; (ii) can handle larger step sizes effectively; and (iii) performs well even in the presence of higher levels of measurement noise. As the first illustration of learning SDEs, we examine a two-dimensional system that has been studied for computing invariant distributions from data [29]. In that research, the Euler-Maruyama approach is employed to learn the governing functions, and the invariant distribution is computed simultaneously. Here, we use the proposed DynGMA to reconstruct the entire dynamics and then compute the invariant distribution using their decomposition method. The system is formulated as:
@@
We generated $2000\left(\frac{\Delta t}{0.05}\right)$ trajectories for our training data. Each trajectory consists of an initial state uniformly sampled from $[-2,2]\times[-3,3]$, and includes $M=1/\Delta t$ states at equidistant time steps of $\Delta t$. In other words, there are a total of $4\times 10^{4}$ data points in the dataset. We employed an FNN with two hidden layers, each having 128 units, and using tanh activation function to parameterize the drift $f_{\theta}$. Additionally, We use a constant $2$-by-$2$ matrix to parameterize the diffusion $\sigma_{\theta}$. We use Algorithm 1 to approximate the transition density. Here the step size of each sub-Gaussian step is set to $h=0.1$ and the step size of discretization (14) is fixed at $0.05$. The number of steps is determined as $K=\lceil\Delta t/h\rceil$ and $L=\min(h,\Delta t)/0.05$. Specifically, for $\Delta t$ values of $0.05$, $0.1$, and $0.2$, the values of $K$ are $1,1,2$, and the values of $L$ are $1,2,2$. During training, we employed full-batch Adam optimization for $5\times 10^{3}$ epochs, with the learning rate set to exponentially decay with linearly decreasing powers from $10^{-2}$ to $10^{-4}$.
@@
Figure 4: Time evolution of the distribution (Top), the mean and the standard deviation (Bottom) of the learned and exact dynamics for the two-dimensional system.
@@
After completing the training process, we compute the invariant distributions and the generalized potential of the learned dynamics using the decomposition method [29], of which details are present in B. Subsequently, we plot the drift functions and the generalized potential of the learned and the exact dynamics in Fig. 3. Here, the reference solution is derived by numerically solving the FPK equation using the finite difference method. The results demonstrate agreements between both solutions when the step size $\Delta t$ is small. However, the Euler-Maruyama approach exhibits limitations in learning accurate dynamics and generalized potentials for larger step sizes. In contrast, DynGMA consistently provides accurate results, as evidenced in the third column of Fig. 3.
Subsequently, we uniformly sample $2\times 10^{3}$ initial points from the domain $[-2,2]\times[-3,3]$. Next, in the top portion of Fig. 4, we present the distribution of these initial states and their subsequent evolution, following both the exact dynamics and the dynamics learned by DynGMA when $\Delta t=0.2$. In the bottom portion of Fig. 4, we display the mean and standard deviation of both the exact trajectories and learned trajectories for $\Delta t=0.05,0.1,0.2$. The results demonstrate that the SDE learned using DynGMA more accurately captures the true dynamics.
@@
Figure 5: Errors in the learned diffusion coefficients and the generalized potential of the learned dynamics for multi-step loss
when using data with measurement noise for the two-dimensional system (19). The results are obtained by taking the mean of $5$ independent experiments, and the black error bar represents one standard deviation.
Next, we show that multi-step training using DynGMA improves robustness in the presence of measurement noise in the system observations. We generate $4000$ trajectories and sample $41$ states at equidistant time steps with $\Delta t=0.025$. Each trajectory point is multiplied by $1+\delta$ to generate the training dataset, where $\delta$ follows a uniform distribution on $[-\text{Noise},\text{Noise}]$, and the intensity Noise is set to $0.01$, $0.02$, and $0.04$. We optimize the multi-step loss (18), where $\Gamma=\{j\}$ and $j$ takes values $1,2,4,8$. As the Euler-Maruyama approach is only valid for one step, we directly expand its step size, take
@@
and substitute it into loss (18). Note that excessive measurement noise makes it challenging to distinguish between measurement noise and the true randomness induced by diffusion. Thus, we record the error of the learned diffusion and the learned potential in Fig. 5. It depicts that the error increases as the noise intensity increases when employing small steps ($j=1,2$) for both methods. With multi-step training involving more steps, the discrepancy in error levels among the results obtained using DynGMA under different noise intensities diminishes. However, augmenting the number of steps in the Euler-Maruyama approach results in larger time step sizes, consequently resulting in notable numerical errors. In contrast, DynGMA is capable of handling data with larger step sizes, allowing it to deal with measurement noise by combining multi-step training.
Figure 6: (Left and Middle) convergence of error of DynGMA and the Euler-Maruyama approach with respect to time step size $\Delta t$. (Right) convergence of error of DynGMA with respect to parameter $L$.
Finally, to numerically verify the accuracy of the Gaussian density approximation demonstrated in Theorem 3.2, 3.3 and 3.4, we test the convergence of the errors incurred by Algorithm 1 (with $K=1$ and $L=2$), Algorithm 2 (with $K=1$), and the Euler-Maruyama approach across varying data step sizes ($\Delta t=0.05,0.1,0.2,0.4$), while retaining the same data generation parameters and network settings. The left and middle panels of Fig. 6 illustrate the convergence of errors for DynGMA and the Euler-Maruyama approach with respect to $\Delta t$. It is observed that the errors of all methods decrease as $\Delta t$ decreases. Eventually, both DynGMA methods reach a plateau where the rate of error reduction levels off. In addition to the discretization error induced by density approximation, the learning errors of neural networks—encompassing generalization, optimization, and approximation errors—can also influence the final outcome. Therefore, to calculate the convergence order, we construct linear regression using only the errors for $\Delta t=0.1,0.2,0.4$ where numerical error accounts for the dominant term. The black solid line here represents the result of regression. DynGMA with Algorithms 1 and 2 consistently demonstrate a higher convergence order. The right panel of Fig. 6 shows the convergence of error for DynGMA using Algorithm 1 with respect to parameter $L$. Increasing $L$ from $1$ to $2$ results in a notable decrease in error. However, further increments in $L$ do not yield significant improvement as learning error may dominated. As our theoretical analysis focuses on the error of the density approximation, it is not able to directly quantify the error between the true governing function and its approximation. We would like to further investigate this problem in future research.
@@ 4.3 Learning SDEs with state-dependent diffusions
Subsequently, we consider the $3$d Lorenz system [32, 44], which is a simplified model for atmospheric dynamics and is widely investigated for learning dynamics [29, 48]. To showcase the effectiveness of the proposed method in learning SDEs with state-dependent diffusions, we investigate the Lorenz system in the presence of multiplicative noise [9], formulated as:
@@
We simulate $N=10^{4}$ trajectories for the training dataset. For each trajectory, the initial state is sampled from the uniform distribution on the domain $[-25,25]\times[-30,30]\times[-10,60]$, and we collect $M=200$ pairs of data points at the time slices $\{t_{j}=0.01*j:\ j=0,1,\cdots,M-1\}$. In total, there are $2\times 10^{6}$ pairs of points. Here, Algorithm 2 with $h=0.005$, corresponding to $K=2$, is used to obtain the approximation of the transition density. And the results are compared with those obtained using the Euler-Maruyama and Gaussian Cubature methods (employing the Euler method with a step size of $0.005$). In all methods, the drift $f_{\theta}$ is parameterized in the same manner as described in Section 4.2, and the diffusion $\sigma_{\theta}$ is parameterized according to (16), where the used FNN has one hidden layer with $50$ units and tanh activations. The optimization process employs the Adam optimization algorithm, with a chosen batch size of $10^{5}$. The learning rate is set to $10^{-3}$ for DynGMA and the Euler-Maruyama and $10^{-4}$ for the Gaussian Cubature approach. Results are collected after $10^{5}$ parameter updates.
@@
Moreover, we compute the errors $e_{f}$, $e_{\sigma}$, and $e_{p}$, evaluated at $10^{5}$ points uniformly sampled in $[-25,25]\times[-30,30]\times[-10,60]$. The results are presented in Table 3 for various methods. It is shown that our proposed DynGMA outperforms other methods in terms of accuracy. Interestingly, despite the Gaussian Cubature having a discrete step size smaller than that of the Euler-Maruyama approach, it fails to yield a substantial improvement. It is noted that the Gaussian Cubature method involves two steps, and an excessive number of steps can potentially lead to non-positive covariance, resulting in the failure of the training process. Currently, the application of the Gaussian Cubature method in machine learning is limited to time series prediction using SDE as black-box [42]. Its utilization for learning interpretable drift and diffusion functions in SDE models has not been explored. Further research might be required to investigate the parameterization and discretization of this method for such tasks.
@@ 4.4 Learning high-dimensional SDEs
We next validate the performance of the DynGMA in learning a $10$d biological model combined by the epithelial-mesenchymal transition (EMT) network and the metastasis network. Researchers can investigate the correlation between EMT and the metastasis of cancer via this biological model [19] and have proposed several works to compute its invariant distribution [26, 29]. The Euler-Maruyama loss has been employed to learn its stochastic dynamics and invariant distributions from data [29]. Herein, we use a similar experimental setup but a larger time step size.
Specifically, the training data is obtained by solving the governing SDE with constant diffusion, of which the governing function is given in A. We first sample $10^{5}$ initial states from the uniform distribution on the domain $\Omega_{1}\times\cdots\times\Omega_{10}$ where $\Omega_{i}=[0,2]$ for $i\neq 7$ and $\Omega_{7}=[0,6]$, then simulate trajectories and collect $36$ pairs of data points at the times $\{t_{j},t_{j}+\Delta t\}$, where $t_{j}=2,2.5,\cdots,19.5$. In total, there are $7.2\times 10^{6}$ data points. The step size $\Delta t$ is set as $0.005$ in previous work [29] and here we configure $\Delta t$ as $0.04$ or $0.08$. Due to considerations of computational efficiency and resources, we apply the Adam optimization with a batch size of $10^{5}$ for $10^{5}$ iterations to train the model, and we use Algorithm 1 with $K=1$ and $L=2$, i.e., Gaussian density approximation with (14). It is noted that in high-dimensional scenarios, DynGMA with larger $K$ would significantly increase computational complexity. Therefore, herein we solely employ the proposed Gaussian approximation. The chosen model architecture and other hyperparameters remain consistent with those outlined in Section 4.2.
@@ 4.5 Learning effective SDEs from Gillespie’s stochastic simulations data with variable time step-sizes
In the subsequent sections, we validate the effectiveness of the proposed DynGMA in handling data with variable, and even uncontrollable, time step-sizes. We consider learning effective SDEs from trajectories generated by Gillespie’s stochastic simulation algorithm (SSA), where an exact SDE does not exist. SSA generates sample sequences of continuous-time Markov chains, where individual events occur with random timing. The time steps between reaction events follow an exponential distribution, resulting in variable and uncontrollable step sizes. Specifically, We consider the Susceptible-Infectious-Recovered (SIR) model, an epidemiological compartment model used to simulate the spread of infectious diseases in a population. In this model, individuals can transition between three states: Susceptible (S), Infectious (I), and Recovered (R). For a total population size of $N$, we denote the counts of susceptible, infected, and recovered individuals as $n_{0}$, $n_{1}$, and $n_{2}$, respectively. Their corresponding concentrations are represented by $y_{0}$, $y_{1}$, and $y_{2}$. At each time step of Gillespie’s SSA, the rates $r_{1}=4k_{1}y_{0}y_{1}$, $r_{2}=k_{2}y_{1}$, and $r_{3}=k_{3}y_{2}$ are computed, and subsequently, $y_{1}$ and $y_{2}$ are updated accordingly. The mass balance then determines that the concentration of susceptible individuals is given by $y_{0}=1-y_{1}-y_{2}$.
@@
In Gillespie’s SSA, the parameters $t_{max}$ and $\Delta t$ are predetermined, and the data points are evaluated at ${t_{0},t_{1},\cdots,t_{max}}$, where $\Delta t_{i}=t_{i}-t_{i-1}\approx\Delta t$. In this experiment, the model parameter settings for the SIR model are as follows: $k_{1}=1$, $k_{2}=1$, $k_{3}=0$ and $N=1024$. We set $t_{max}=1$ and choose various values for $\Delta t$ to generate $10^{5}$ trajectories for each training dataset and $2.5\times 10^{4}$ trajectories for each test dataset. The data step size is a random variable and is illustrated in Fig. 9 for different settings of $\Delta t$. There exist some uncontrollably large steps $\Delta t_{i}$ for all settings of $\Delta t$. We use one hidden layer in FNN with $64$ units and tanh activations to represent drift $f_{\theta}$, and parameterize diffusion $\sigma_{\theta}$ according to (16). Adam optimization with the learning rate of $10^{-3}$ is used to update parameters for $2\times 10^{4}$ epochs. Under these settings, we carry out Algorithm 2 (if $\Delta t=0.2$, then $K=2$; otherwise, $K=1$) to obtain the approximation and optimize the loss function (17) to train the networks.
@@
Next, we consider the SIRS model by setting $k_{3}=0.5$, where recovered individuals have a probability of being reinfected and richer dynamic behaviors are exhibited. We generate the data for the SIRS model in the same manner and use the same settings to train networks. The predictions are shown in Fig. 12. It is shown that the learned network model captures stochastic dynamics perfectly and is capable of reconstructing the correct oscillations and steady state.